### v1

# Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter

Abstract
This work comprises the calculations of magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in the presence of hot and dense hadronic matter. Using the chiral $SU(3)$ quark mean field model, we have taken into account the impact on in-medium scalar meson fields to investigate the density effects on the in-medium baryon masses and their constituent quarks. In light of chiral constituent quark model $\chi$CQM, we have calculated the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances and scrutinized the explicit contributions coming from the valence quarks, sea quarks and the orbital angular momentum of sea quarks. The effective baryonic magnetic moments have been investigated for the varying values of isospin asymmetry and strangeness fraction.
I Introduction
In quantum chromodynamics (QCD), quarks and gluons are treated as fundamental degrees of freedom which form bound states known as hadrons (baryons($qqq$) and mesons($q\bar{q}$)). These fundamental degrees of freedom undergo dynamical changes due to strong interactions in dense nuclear medium, which in turn cause in-medium modifications in the properties of hadrons. Confinement makes it challenging to understand QCD at lower energy scales. Scrutinizing the internal structure and the properties of light hadrons allow us to unravel the physics of strong interactions at low energies. Apart from heavy-ion collision experimental studies on chiral symmetry restoration at high temperature thirtyseven , and the de-confinement of hadrons to quark-gluon plasma (QGP) thirtyeight , it has become a primary interest for experiments such as PANDA at FAIR thirtythree ; thirtyfour , BESIII at BEPCII thirtysix ; thirtytwo , HADES at GSI thirtyfive etc., to explore the electromagnetic properties of hadrons in dense matter. Electromagnetic form factors, being key observable to investigate the electric and magnetic moment distributions, are of theoretical interest to explore other primary hadronic properties like magnetic moment and charged radii to probe deep inside of hadrons.
Magnetic moment of hadrons, in free space, play a very important role in investigating their internal structure at a sub-nuclear level sixtyfour . Theoretically, the magnetic moments of octet as well as decuplet baryons have been extensively studied in free space one ; two ; three ; four ; five ; six ; seven ; eight ; nine . Isgur framed his model with one-gluon exchange, in order to account for internal structure of hadrons isgur1 ; isgur2 ; isgur3 ; isgur4 . The constituent quark model (CQM) propounded that the baryonic magnetic moments can be calculated as the summation of the individual magnetic moments of its constituent quarks ten ; eleven . It was Glozman who incorporated the Goldstone boson exchange (GBE) interaction in the low energy regime and the concept of one-gluon exchange was substituted by vector meson exchange interactions, which in fact led to better fits for baryon spectra glozman . It was not until the development of chiral constituent quark model ($\chi\mathrm{CQM}$) that the pseudo-scalar and scalar meson exchanges were considered which was later extended to the inclusion of vector meson exchange dai . Hence, $\chi\mathrm{CQM}$ was formed by extending the quark model and further incorporating the chiral symmetry and corresponding spontaneous symmetry breaking into it twelve . The $\chi\mathrm{CQM}$ model provides satisfactory explanation to the “Proton spin puzzle”, which was an outcome of the final data taking at EMC, and other hadron properties ten ; thirteen ; fourteen ; fifteen ; sixteen ; seventeen ; eighteen ; nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; twentyfour ; twentyfive ; twentysix ; twentyseven . In previous studies, this model has been used for calculating the baryonic magnetic moments, by taking into account the sea quark polarizations and their orbital angular momentum contribution through the Cheng-Li mechanism twentyeight ; twentynine ; thirty . In the $\chi$CQM formalism, the magnetic moments of low-lying ${\frac{1}{2}}^{-}$ $N^{*}$ resonances thirtyone , ${\frac{1}{2}}^{-}$ octet baryon resonances thirtyone and low-lying ${\frac{1}{2}}^{-}$ $\Lambda$ resonances A_Torres have been calculated in earlier studies. Apart from $\chi$CQM, the magnetic moments of $N^{*}$ resonances have been computed by three quark Hamiltonian approach and the results were not consistent with lattice QCD values with very large variations hamiltonian_approach . As an extension to chiral perturbation theory, the Chiral unitary model was also used in ref. cum_hyodo to calculate the magnetic moments of $N^{*}(1535)$ resonances.
As mentioned earlier, the study of magnetic moments of baryons and the medium dependent effects on the electromagnetic properties is an area of active research and stands relevant to some of the experiments. In order to study the properties of nuclear matter, various theoretical models like the MIT bag model thirtynine , linear sigma model Petro , non-linear sigma model fortyfive , Zimanyi and Moszkowski model seventyone , cloudy bag model fortythree , Nambu-Jona-Lasinio (NJL) model fortytwo , light-front quark model Schlumpf1993 , parity doublet model paritydoublet1 ; paritydoublet2 etc., have been developed. The cloudy bag model attained increased accuracy over MIT bag model in explaining QCD effects by incorporating the presence of pion clouds at the core of a baryon. With further development, models like the NJL model successfully explained low energy baryonic dynamics by considering chiral quarks as constituents of hadrons, through the inclusion of the interaction between the constituent quarks njlsuccess . Few years later, the quark meson coupling (QMC) model was constituted and it facilitated a deeper understanding of the in-medium baryon properties fortyfour . The QMC model, formulated in lines of cloudy bag and NJL model, has previously been deployed to study the impact of medium on the magnetic moments of octet baryons and has produced results which are in good agreement with experimental data seventyfour . The chiral $SU(3)$ quark mean field model (CQMF) fortyfour ; fortyfive ; fortysix ; fortyseven is an extended version of the quark mean field model fortyeight , which is based on the QMC model approach. In the CQMF model, the constituent quarks are confined into baryons by a confining potential based on mean field approximation using classical expectation values instead of quantum field operators fortynine . In the context of medium modifications of electromagnetic properties of hadrons, the studies in ref. kaon show the in-medium effects on $K^{+}$ meson properties in light-front quark meson coupling model (LF-QMC) in symmetric nuclear matter. With the rise in medium density, the kaon electromagnetic form factors decrease faster than in vacuum, while the kaon electromagnetic charge radii grow in size. By studying pion and kaon meson structures in NJL model the authors in ref. njlmeson have obtained results consistent with that in ref. kaon . For octet baryons in light of the QMC model effoctet , it has been shown that the in-medium electric form factors are dampened and the magnetic form factors are enhanced, and this finding is consistent with the in-medium variations of the bound proton electromagnetic form factors observed at Jefferson Lab jlab . The study of effective magnetic moments of baryons was also extended to calculate the low-lying charm, and the low-lying bottom baryons with the aim of understanding the roles of the heavy and light quarks in symmetric nuclear matter tsushima1 ; tsushima2 . While a comparative study on octet baryons with modified quark meson coupling model (MQMC) mqmc , suggests that the observed in-medium baryonic magnetic moments incur drastic variations compared to that in QMC framework. In CQMF formalism, the medium modifications of magnetic moments of octet fiftythree , decuplet fiftyfour and charmed Dutt:2024lui baryons have been studied in hot and dense hadronic matter. The CQMF model has recently been used to study valence quark distribution and transverse momentum dependent distributions of valence quarks of pions Puhan:2024xdq ; Kaur:2024wze .
Presence of strangeness fraction ($f_{s}$) is a principal aspect to understand the departure of hadronic properties of baryons in strange hadronic matter comprising of hyperons from that in a non-strange medium. The in-medium variations of magnetic moments of octet and decuplet baryons in strange medium has shown stark deviation, as studied in fiftythree ; fiftyfour ; sixtytwo , for baryons of higher strangeness quantum number. In the present work, we have focused our study on the in-medium effects on the magnetic moments of the low-lying $N^{*}$ resonances and $\frac{1}{2}^{-}$ octet baryon resonances in CQMF model. It is exciting to study these resonances, as their detection connects hadron models to QCD properties in low and intermediate energy regimes. Photoproduction has been a key experimental tool involved in detecting the low-lying states of baryon resonances as shown in Refs. fifty ; fiftyone ; fiftytwo . The large decay widths of these excited baryon resonances, makes it challenging to detect them in experiments. In our current work, we have incorporated $\chi\mathrm{CQM}$ to calculate the properties of the baryon resonances as in thirtyone and have visualized them as a system of three valence quarks surrounded by the meson cloud. We have calculated the in-medium magnetic moments of the ${\frac{1}{2}}^{-}$ baryon resonances at finite temperature and density of hadronic medium through the medium modification of the corresponding constituent quark masses and the baryon masses. Within the CQMF framework, the effective masses of quarks ($m^{*}_{q}$) are calculated through the medium modified values of the scalar fields $\sigma$ and $\zeta$ and scalar iso-vector field $\delta$. The confined quarks inside the baryons, participate in the exchange interaction with these scalar meson fields as well as vector meson fields $\omega$, $\rho$ and $\phi$.
This manuscript has been organized as follows: In section II the CQMF model used to compute the in-medium quark masses and effective baryon masses, at finite temperature and density of the medium is presented. In section III we have calculated the contributions of valence quarks, sea quarks and orbital angular momentum of sea quarks towards the total magnetic moments of baryons, in light of $\chi\mathrm{CQM}$ model. Section IV is dedicated to a detailed discussion on the results obtained from numerical calculations. We devote section V to present the summary of the current work.
II Chiral Quark Mean Field Model
In this section, we discuss briefly the framework, CQMF model, to calculate the in-medium constituent quark and baryon masses. As discussed earlier, in the CQMF model the quarks and mesons are treated as fundamental degrees of freedom, and the quarks are assumed to be confined inside baryons by an effective confining potential. The constituent quarks and mesons obtain their masses through the inclusion of properties such as spontaneous and explicit (for pseudoscalar mesons) symmetry breaking fortyfour . In the CQMF model, the effective Lagrangian density accounting for various interaction terms is given as fortyfour ; fortyfive ; fiftythree
${\cal L}_{{\rm eff}}\,=\,{\cal L}_{q0}\,+\,{\cal L}_{qm}\,+\,{\cal L}_{VV}\,+\,{\cal L}_{\Sigma\Sigma}\,+\,{\cal L}_{\chi SB}\,+\,{\cal L}_{c},$ (1)
where $\,{\cal L}_{q0}=\bar{q}i\gamma^{\mu}\partial_{\mu}q\,$ is the kinetic term for the free part of massless quarks and ${\cal L}_{qm}\,$ is the quark-meson interaction term that remains invariant under chiral $SU(3)$ transformations. ${\cal L}_{qm}\,$ can be written as fortyfour ; fortysix ; fiftythree ; fiftyfour
$\mathcal{L}_{qm}$ $=g_{\mathrm{s}}\left(\bar{\Psi}_{L}M\Psi_{R}+\bar{\Psi}_{R}M^{\dagger}\Psi_{L}\right)-g_{v}\left(\bar{\Psi}_{L}\gamma^{\mu}l_{\mu}\Psi_{L}+\bar{\Psi}_{R}\gamma^{\mu}r_{\mu}\Psi_{R}\right)$ (2)
$=\frac{g_{\mathrm{s}}}{\sqrt{2}}\bar{\Psi}\left(\sum_{a=0}^{8}s_{a}\lambda_{a}+\mathrm{i}\gamma^{5}\sum_{a=0}^{8}p_{a}\lambda_{a}\right)\Psi-\frac{g_{v}}{2\sqrt{2}}\bar{\Psi}\left(\gamma^{\mu}\sum_{a=0}^{8}v_{\mu}^{a}\lambda_{a}-\gamma^{\mu}\gamma^{5}\sum_{a=0}^{8}a_{\mu}^{a}\lambda_{a}\right)\Psi.$
In eq. 2, $\Psi=$ $\begin{pmatrix}u\\ d\\ s\end{pmatrix},$ and $g_{\mathrm{s}}$ and $g_{v}$ are the respective scalar and vector coupling constants. The relations between various quark-meson coupling constants are given as fortysix
$\frac{g_{s}}{\sqrt{2}}=$ $g_{\delta}^{u}=-g_{\delta}^{d}=g_{\sigma}^{u}=g_{\sigma}^{d}=\frac{1}{\sqrt{2}}g_{\zeta}^{s},~{}~{}~{}~{}~{}g_{\delta}^{s}=g_{\sigma}^{s}=g_{\zeta}^{u}=g_{\zeta}^{d}=0\,,$ (3)
$\frac{g_{v}}{2\sqrt{2}}=g_{\rho}^{u}=-g_{\rho}^{d}=g_{\omega}^{u}=g_{\omega}^{d},~{}~{}~{}~{}~{}~{}~{}g_{\omega}^{s}=g_{\rho}^{s}=0.$ (4)
It is important to note that spin-0 scalar ($\Sigma$) and pseudoscalar ($\Pi$) meson nonets can be expressed as fortyfour ; fortysix
$M\left(M^{\dagger}\right)=\Sigma\pm i\Pi=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}\left(s^{a}\pm i\pi^{a}\right)\lambda^{a},$ (5)
where $s^{a}$ and $\pi^{a}$ represent the nonets of scalar and pseudoscalar mesons, respectively, while $\lambda^{a}$ are Gell-Mann matrices with $\lambda^{0}=\sqrt{\frac{2}{3}}I$. The spin-1 mesons can also be defined in a similar way as
$l_{\mu}\left(r_{\mu}\right)=\frac{1}{2}\left(V_{\mu}\pm A_{\mu}\right)=\frac{1}{2\sqrt{2}}\sum_{a=0}^{8}\left(v_{\mu}^{a}\pm a_{\mu}^{a}\right)\lambda^{a}.$ (6)
In the above expression, $v_{\mu}^{a}$ and $a_{\mu}^{a}$ are nonets of vector and pseudovector mesons, respectively. The physical states of scalar and vector meson nonets are given by
$\Sigma=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}s_{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\sigma+\delta^{0}\right)&\delta^{+}&\kappa^{*+}\\ \delta^{-}&\frac{1}{\sqrt{2}}\left(\sigma-\delta^{0}\right)&\kappa^{*0}\\ \kappa^{*-}&\bar{\kappa}^{*0}&\zeta\end{array}\right),$ (10)
and
$V_{\mu}=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}v_{\mu}^{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\omega_{\mu}+\rho_{\mu}^{0}\right)&\rho_{\mu}^{+}&K_{\mu}^{*+}\\ \rho_{\mu}^{-}&\frac{1}{\sqrt{2}}\left(\omega_{\mu}-\rho_{\mu}^{0}\right)&K_{\mu}^{*0}\\ K_{\mu}^{*-}&\bar{K}_{\mu}^{*0}&\phi_{\mu}\end{array}\right),$ (14)
respectively.
The term, ${\cal L}_{VV}$ in eq. 1 represents the chiral-invariant vector meson self interaction terms. Within the framework of mean field approximation ${\cal L}_{VV}$ is given by fortyfour ; fiftythree ; fiftyfour
${\cal L}_{VV}=\frac{1}{2}\,\frac{\chi^{2}}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}+m_{\phi}^{2}\phi^{2}\right)+g_{4}\left(\omega^{4}+6\omega^{2}\rho^{2}+\rho^{4}+2\phi^{4}\right).$ (15)
For the scalar mesons, their self interaction terms are written as (fourth term in eq. 1)
${\cal L}_{\Sigma\Sigma}=$ $-\frac{1}{2}\,k_{0}\chi^{2}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)+k_{1}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)^{2}+k_{2}\left(\frac{\sigma^{4}}{2}+\frac{\delta^{4}}{2}+3\sigma^{2}\delta^{2}+\zeta^{4}\right)$
$+k_{3}\chi\left(\sigma^{2}-\delta^{2}\right)\zeta-k_{4}\chi^{4}-\frac{1}{4}\chi^{4}{\rm ln}\frac{\chi^{4}}{\chi_{0}^{4}}+\frac{\xi}{3}\chi^{4}{\rm ln}\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\left(\frac{\chi^{3}}{\chi_{0}^{3}}\right)\right).$ (16)
The parameters $k_{0},k_{1},k_{2},k_{3}$ and $k_{4}$ appearing in the above Lagrangian density are calculated using masses of $\pi$ and $K$ as well as the average mass of $\eta$ and $\eta^{{}^{\prime}}$ mesons. The vacuum expectation values of scalar meson fields $\sigma$ and $\zeta$, i.e. $\sigma_{0}$ and $\zeta_{0}$, can be expressed in terms of pion and kaon leptonic decay constants $f_{\pi}$ and $f_{K}$, through the relations
$\sigma_{0}=-f_{\pi}~{}~{}{\rm and}~{}~{}~{}~{}\zeta_{0}=\frac{1}{\sqrt{2}}\left(f_{\pi}-2f_{K}\right),$ (17)
where $f_{\pi}=92.8$ MeV and $f_{K}=115$ MeV. The vacuum value of dilaton field $\chi_{0}=254.6$ MeV and the coupling constant $g_{4}=37.4$, obtained from fit reasonable effective nucleon mass fortyfour ; fiftythree ; fiftyfour . The last three terms of eq. 16 are introduced in the model to incorporate the trace anomaly property of QCD, which leads to the trace of energy momentum tensor being proportional to fourth power of dilaton field $\chi$ fortynine . The parameter $\xi$, originating from the logarithmic term used in eq. 16, is usually determined from QCD $\beta$-function at one loop level for three colors and flavors fortynine . The fifth term of the effective Lagrangian density of eq. 1 is the explicit chiral symmetry breaking term, ${\cal L}_{\chi SB}$, and is expressed as
${\cal L}_{\chi SB}=\frac{\chi^{2}}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}\,m_{K}^{2}f_{K}-\frac{m_{\pi}^{2}}{\sqrt{2}}f_{\pi}\right)\zeta\right].$ (18)
The above expression is introduced in chiral effective models to account for the non-vanishing pseudoscalar meson masses and it satisfies the partial conserved axial-vector current relations for $\pi$ and $K$ mesons fortyfour ; fortysix ; fortyseven . The confinement of constituent quarks inside the baryons, due to the presence of a confining scalar-vector potential $\chi_{c}$, is introduced through the last term in the Lagrangian density, in eq. 1, and is written as
${\cal L}_{c}=-\bar{\psi}\chi_{c}\psi,$ (19)
where the confining potential $\chi_{c}$ is given by
$\chi_{c}(r)=\frac{1}{4}k_{c}\,r^{2}(1+\gamma^{0})\,.$ (20)
In eq. 20 above the coupling constant $k_{c}$ is taken to be $98\,\text{MeV}.\text{fm}^{-2}$.
In order to investigate the properties of asymmetric strange matter at finite temperature and density, we are using mean field approximation fortyfour . Under the influence of meson mean field, we express the Dirac equation for the quark field $\Psi_{qi}$ as
$\left[-i\vec{\alpha}\cdot\vec{\nabla}+\chi_{c}(r)+\beta m_{q}^{*}\right]\Psi_{qi}=e_{q}^{*}\Psi_{qi}.$ (21)
Here the quark $q$ ($q=u,d,s$) is in a baryon of type $i$ ($i=N,\Lambda,\Sigma,\Xi$)  and $\vec{\alpha}$ , $\beta$  are the usual Dirac matrices. The effective quark mass $m_{q}^{*}$ is defined as
$m_{q}^{*}=-g_{\sigma}^{q}\sigma-g_{\zeta}^{q}\zeta-g_{\delta}^{q}I^{3q}\delta+m_{q0},$ (22)
At zero baryon density, the masses of constituent quarks $u$, $d$, $s$ are expressed in terms of vacuum expectation values $\sigma_{0}$ and $\zeta_{0}$ of meson scalar fields $\sigma$ and $\zeta$, respectively. For the light $u$ and $d$ quarks, we have
$m_{u}=m_{d}=-g_{\sigma}^{q}\sigma_{0}=-\frac{g_{s}}{\sqrt{2}}\sigma_{0}.$ (23)
The values of coupling constant $g_{s}$ and confining parameter $k_{c}$ are fitted to obtain binding energy of $-16.0$ MeV at nuclear saturation density, $\rho_{0}=0.16$ $\text{fm}^{-3}$. This fit provides the value of effective mass of light quarks to be equal to $256$ MeV. Here $m_{q0}$ is zero for non-strange $u$ and $d$ quarks, whereas for strange $s$ quark $m_{q0}=\Delta m=77$ MeV. The term $\Delta m$ is introduced through the Lagrangian density ${\cal L}_{\Delta m}=-\Delta m\bar{\psi}m_{1}\psi$, where $S_{1}=\frac{1}{3}\left(1-\lambda_{8}\sqrt{3}\right)$. In eq. 21, $e_{q}^{*}$ in the Dirac equation is the effective energy of quarks and is expressed as fortyfour
$e_{q}^{*}=e_{q}-g_{\omega}^{q}\omega-g_{\rho}^{q}I^{3q}\rho-g_{\phi}^{q}\phi.$ (24)
The in-medium mass of a baryon is expressed in terms of its effective energy $E^{*}_{i}$ and spurious center of mass momentum $p_{icm}$ as in Barik1985 ; Barik2013 ,
$M^{*}_{i}=\sqrt{E^{*2}_{i}-<p^{*2}_{icm}>}.$ (25)
Further the in-medium energies of $i^{th}$ baryon can be related to the in-medium energies of quarks and number of q type constituent quarks $n_{qi}$ in $i^{th}$ baryon by the expression
$E^{*}_{i}=\sum_{q}n_{qi}e_{q}^{*}+E_{ispin}.$ (26)
Here, $E_{ispin}$ is a correction term in the effective energy of a baryon that accounts for spin-spin interaction between constituent quarks. The $E_{ispin}$ is determined so as to obtain the correct baryon masses at zero density. Finally, the thermodynamic potential $\Omega$ describing the isospin asymmetric strange hadronic matter at finite temperature and baryon density is written as fortyfour
$\Omega=-\frac{k_{B}T}{(2\pi)^{3}}\sum_{i}\gamma_{i}\int_{0}^{\infty}d^{3}k\left\{\ln\left(1+e^{-\left[E_{i}^{*}(k)-\nu_{i}^{*}\right]/k_{B}T}\right)+\ln\left(1+e^{-\left[E_{i}^{*}(k)+\nu_{i}^{*}\right]/k_{B}T}\right)\right\}-\mathcal{L}_{M}-\mathcal{V}_{\text{vac }},$ (27)
where the mesonic Lagrangian is $\mathcal{L}_{M}=\mathcal{L}_{VV}+\mathcal{L}_{\Sigma\Sigma}+\mathcal{L}_{\chi SB}$ and we have $E^{*}_{i}(k)=\sqrt{M_{i}^{*2}+k^{2}}$. In the above equation, the degeneracy factor is $\gamma_{i}=2$ for baryons and the summation is over hyperons and nucleons. In eq. 27, the effective chemical potential $\nu_{i}^{*}$ is expressed in terms of the usual baryon chemical potential $\mu_{i}$ as fortyfour ; fiftythree ; fiftyfour ; sixtytwo
$\nu_{i}^{*}=\mu_{i}-g_{\phi}^{i}\phi-g_{\omega}^{i}\omega-g_{\rho}^{i}I^{3i}\rho.$ (28)
From eq. 27 the equations of motion of scalar and vector meson fields are calculated by minimizing the effective thermodynamic potential as (see APPENDIX)
$\frac{\partial\Omega}{\partial\sigma}=\frac{\partial\Omega}{\partial\zeta}=\frac{\partial\Omega}{\partial\chi}=\frac{\partial\Omega}{\partial\delta}=\frac{\partial\Omega}{\partial\omega}=\frac{\partial\Omega}{\partial\rho}=\frac{\partial\Omega}{\partial\phi}=0.$ (29)
Equation 29 is solved for finite values of temperature and density of the medium. In order to account for isospin asymmetry in the hadronic medium a parameter $i_{a}=-\frac{\Sigma_{i}I_{3i}\rho_{i}}{\rho_{B}}$ is introduced, where $I_{3i}$ represents the third component of isopin and $\rho_{B}$ is baryonic density of the medium fiftyfour . To take into account hyperons in the medium, we define strangeness fraction $f_{s}=\frac{\Sigma_{i}\left|s_{i}\right|\rho_{i}}{\rho_{B}}$, here $|s_{i}|$ denotes the number of $s$ quarks in $i^{th}$ baryon.
We have eq. 25 which addresses our aim to compute the effective baryon masses, by taking into account the effective constituent quark masses, baryon density and the spurious center of mass motion. In the upcoming section we shift the trajectory of our discussion towards the calculation of effective baryonic magnetic moments, by incorporating the medium modified constituent quark masses and the effective baryon masses.
III Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
In the present section, methodology used to calculate the magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances has been discussed. The formalism of chiral constituent quark model ($\chi\mathrm{CQM}$) thirtyone , has been used to calculate the magnetic moment of the baryons through the following fluctuation process eight ; fiftyfive
$q^{\uparrow\downarrow}\rightarrow\mathrm{GB}+q^{\prime\downarrow\uparrow}\rightarrow{\left(q\overline{q}^{\prime}\right)+q^{\prime\downarrow\uparrow}},$ (30)
where GB represents the emitted Goldstone boson and $q\overline{q}^{\prime}+q^{\prime}$ constitute the sea quarks nineteen ; twenty ; twentyone ; twentytwo ; twentythree . The $\chi\mathrm{CQM}$ model incorporates the concept of confinement and chiral symmetry breaking twelve ; fiftysix , where the massless quarks acquire mass through spontaneous breaking of chiral symmetry twelve . In this approach we consider the emission of a GB, which further splits into a $q\overline{q}^{\prime}$ pair. The coupling between the quark and GBs is described by the effective interaction Lagrangian,
$\mathcal{L}=g_{8}\overline{\mathbf{q}}\left(\Phi+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}I\right)\mathbf{q}=g_{8}\overline{\mathbf{q}}\left(\Phi^{\prime}\right)\mathbf{q}.$ (31)
This quark-GB coupling represented in this Lagrangian is weak enough to consider the fluctuation in eq. 30 as a small perturbation and the higher order fluctuations can therefore be neglected. The process of emission of GB from the quark gives rise to the probability of a spin-flip of the quark. In eq. 31, $\zeta^{\prime}=g_{1}/g_{8},$ where $g_{1}$ and $g_{8}$ denote the coupling constants for the singlet and octet GBs, respectively and $I$ is a $3\times 3$ identity matrix. The matrix of the GBs can be written as
$\Phi^{\prime}=\left(\begin{array}[]{ccc}\frac{\pi^{0}}{\sqrt{2}}+\beta\frac{\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}&\pi^{+}&\alpha K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\beta\frac{\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}&\alpha K^{0}\\ \alpha K^{-}&\alpha\bar{K}^{0}&-\beta\frac{2\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}\end{array}\right)$ (32)
and
$q=\left(\begin{array}[]{l}u\\ d\\ s\end{array}\right).$
The $SU(3)$ chiral symmetry breaking is introduced in $\chi\mathrm{CQM}$ by considering that $m_{s}>m_{u,d}$ and the masses of GBs as non-degenerate $\left(m_{\eta^{\prime}}>m_{K,\eta}>m_{\pi}\right)$ fiftyseven . The parameter $a=\left|g_{8}\right|^{2}$ is denoted as the transition probability of chiral fluctuation $u(d)\rightarrow$ $d(u)+\pi^{+(-)}$, whereas $a\alpha^{2},a\beta^{2}$ and $a\zeta^{\prime 2}$ denote the probabilities of transitions of $u(d)\rightarrow s+K^{-(0)},u(d,s)\rightarrow$ $u(d,s)+\eta$ and $u(d,s)\rightarrow u(d,s)+\eta^{\prime}$, respectively nineteen ; twenty ; twentyone ; twentytwo ; twentythree . By thoroughly analysing the spin and flavor distribution functions of proton, the best fitted set of parameters crucial to our calculations are nineteen ; twentythree ; twentyfive ; twentyeight ; twentynine ; thirty ; thirtyone : $a=0.12,\alpha=\beta=0.45,\eta=-0.15$.
The effective baryonic magnetic moment $\mu^{*}_{B}$ is a summation of the contributions due to spin polarization $\mu^{*S}_{B}$ and the angular momentum polarization $\mu^{*L}_{B}$, expressed as fiftythree ; fiftyfour
$\mu^{*}_{B}=\mu^{*S}_{B}+\mu^{*L}_{B}.$ (33)
Supposedly B ¿ is a SU(6) spin-flavor wave function of a baryon, then we have
$\left\langle B\left|\mu^{*S}_{B}\right|B\right\rangle=\Delta u\mu^{*}_{u}+\Delta d\mu^{*}_{d}+\Delta s\mu^{*}_{s},$ (34)
$\left\langle B\left|\mu^{*L}_{B}\right|B\right\rangle=\Delta u^{(1)}\mu^{*}_{u}+\Delta d^{(1)}\mu^{*}_{d}+\Delta s^{(1)}\mu^{*}_{s}.$ (35)
Here, $\mu^{*}_{q}(q=u,d,s)$ is the effective quark magnetic moment. The spin and orbital angular momentum polarizations are quantitatively comprehended from $\Delta q=q^{\uparrow}-q^{\downarrow}$ and $\Delta q^{(1)}=q^{(1)}-q^{(-1)}$, respectively. The polarizations can be calculated using the number operator $\mathcal{N}$ (=$\mathcal{N}^{\uparrow\downarrow}+\mathcal{N}^{\pm 1}$) as
$\widehat{B}\equiv\langle B|\mathcal{N}|B\rangle.$ (36)
Here $\mathcal{N}^{\uparrow\downarrow}$ and $\mathcal{N}^{(\pm 1)}$ of the number operator $\mathcal{N}$ are expressed as
$\mathcal{N}^{\uparrow\downarrow}=$ $n_{u\uparrow}u^{\uparrow}+n_{u\downarrow}u^{\downarrow}+n_{d\uparrow}d^{\uparrow}+n_{d^{\downarrow}}d^{\downarrow}+n_{s^{\uparrow}}s^{\uparrow}+n_{s\downarrow}s^{\downarrow},$ (37)
$\mathcal{N}^{(\pm 1)}=$ $n_{u^{(1)}}u^{(1)}+n_{u(-1)}u^{(-1)}+n_{d^{(1)}}d^{(1)}+n_{d^{(-1)}}d^{(-1)}+n_{s^{(1)}}s^{(1)}+n_{s^{(-1)}}s^{(-1)},$ (38)
with $n_{q^{\uparrow}}\left(n_{q^{\downarrow}}\right)$ denoting the number of spin up (down) quarks and $n_{q^{(1)}}\left(n_{q^{(-1)}}\right)$ being the number of quarks with the orbital angular momentum projection $m_{L}=1$ $\left(m_{L}=-1\right)$.
Following the Cheng Li mechanism ten , the calculation of total baryonic magnetic moment involves the contribution of valence and sea quark spin polarizations and orbital angular momentum of sea quarks. The spin part in eq. 33, can be obtained from contributions due to valence, sea and orbital angular momentum of sea quarks and is given by
$\mu^{*S}_{B}=\mu^{*S}_{val}+\mu^{*S}_{sea}+\mu^{*S}_{orbit}.$ (39)
The $\mu^{*S}_{val}$, of a baryon, can be derived by calculating the magnetic moments for the physical eigenstates of the orbital angular momentum $L=1$ ($P-$wave) negative-parity resonances thirtyone . For each baryon, the contribution of sea quark spin polarization in spin part of the total magnetic moment is obtained by substituting for each valence quark
$q^{\uparrow(\downarrow)}\rightarrow-P_{[q,GB]q^{\uparrow(\downarrow)}}+\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}.$ (40)
In eq. 40, $P_{[q,GB]}$ is the probability of emission of GBs from a quark $q^{\uparrow(\downarrow)}$ fiftyeight given by
$P(u\rightarrow\mathrm{GB})=P(d\rightarrow\mathrm{GB})=\frac{a}{6}\left(9+6\alpha^{2}+\beta^{2}+2\zeta^{\prime 2}\right),$ (41)
$P(s\rightarrow\mathrm{GB})=\frac{a}{3}\left(6\alpha^{2}+2\beta^{2}+\zeta^{\prime 2}\right),$
and $\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}$ represents the probability of transforming a $q^{\uparrow(\downarrow)}$ quark and it is given by nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; fiftythree ; fiftyfour
$\left|\psi\left(u^{\uparrow(\downarrow)}\right)\right|^{2}=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\uparrow(\downarrow)}+ad^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$ (42)
$\left|\psi\left(d^{\uparrow(\downarrow)}\right)\right|^{2}=au^{\uparrow(\downarrow)}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$
$\left|\psi\left(s^{\uparrow(\downarrow)}\right)\right|^{2}=a\alpha^{2}u^{\uparrow(\downarrow)}+a\alpha^{2}d^{\uparrow(\downarrow)}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\uparrow(\downarrow)}.$
The orbital angular momentum contribution due to sea quarks, is expressed as ten ; fourteen ; twentyeight ; twentynine ; thirty ; thirtyone
$\mu^{*}\left(q^{\uparrow}\rightarrow q^{\prime\downarrow}\right)=\frac{e_{q}}{2m^{*}_{q}}\left\langle l_{q}\right\rangle+\frac{e_{q}-e_{q\prime}}{2M_{\mathrm{GB}}}\left\langle l_{\mathrm{GB}}\right\rangle,$ (43)
where $l_{q}\right>=\frac{M_{\mathrm{GB}}}{M_{\mathrm{GB}}+m^{*}_{q}}$ and $l_{\mathrm{GB}}\right>=\frac{m^{*}_{q}}{M_{\mathrm{GB}}+m^{*}_{q}}$. The quantities $\left(l_{q},l_{\mathrm{GB}}\right)$ and $\left(m^{*}_{q},M_{\mathrm{GB}}\right)$ stands for the orbital angular momenta and masses of quarks and GBs, respectively. The orbital moment of each process in eq. 43, when multiplied by the probability of occurrence of the process, produces the magnetic moment due to all possible transitions for a given valence quark. This can be mathematically expressed as,
${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(u_{\uparrow}\rightarrow u_{\downarrow}\right)\right.$ (44)
$\left.+\mu^{*}\left(u_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(u_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\mu^{*}\left(d_{\uparrow}\rightarrow u_{\downarrow}\right)+\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\right.$
×$\left.\times\mu^{*}\left(d_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(d_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow u_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow d_{\downarrow}\right)\right.$
$\left.+\left(\frac{2}{3}\beta^{2}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(s_{\uparrow}\rightarrow s_{\downarrow}\right)\right].$
The orbital moments of the $u,d$ and $s$ quarks can also be expressed in terms of the $\chi\mathrm{CQM}$ parameters $(a,\alpha,\beta,\zeta^{\prime})$, masses of GBs $\left(M_{\pi},M_{\eta\prime},M_{K},M_{\eta}\right)$ and effective quark masses $\left(m^{*}_{u},m^{*}_{d},m^{*}_{s}\right)$, respectively as thirtyone ; fiftythree ; fiftyfour :
${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\frac{3m_{u}^{*2}}{2M_{\pi}\left(m_{u}^{*}+M_{\pi}\right)}-\frac{\alpha^{2}\left(M_{K}^{2}-3m_{u}^{*2}\right)}{2M_{K}\left(m_{u}^{*}+M_{K}\right)}\right.+\frac{\beta^{2}M_{\eta}}{6\left(m_{u}^{*}+M_{\eta}\right)}$
$\left.+\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{3\left(m_{u}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{u},$
${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{3\left(M_{\pi}^{2}-2m_{d}^{*2}\right)}{4M_{\pi}\left(m^{*}_{d}+M_{\pi}\right)}-\frac{\alpha^{2}M_{K}}{2\left(m_{d}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{12\left(m_{d}^{*}+M_{\eta}\right)}\right.$ (45)
$\left.-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{d}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{d},$
${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{\alpha^{2}\left(M_{K}^{2}-3m_{s}^{*2}\right)}{2M_{K}\left(m_{s}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{3\left(m_{s}^{*}+M_{\eta}\right)}-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{s}^{*}+M_{\eta^{\prime}}\right)}\right]\mu_{s}^{*}.$
In the present study, we have included the effect of relativistic quark confinement and have calculated the mass adjusted effective magnetic moments of its constituent quarks by the expression fiftyfive
$\mu_{d}^{*}=-\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{s}^{*}=-\frac{m_{u}^{*}}{m_{s}^{*}}\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{u}^{*}=-2\mu_{d}^{*}.$ (46)
In the above relations, $\Delta M=M_{B,vac}-M_{B}^{*}$ where $M_{B,vac}$ is the vacuum mass of the given baryon and $M_{B}^{*}$ denotes the in-medium mass of this baryon.
In light of this formalism, we can obtain the spin contribution ($\mu_{B}^{*S}$) towards the baryonic magnetic moment ($\mu^{*}_{B}$). The expressions for spin contribution of n* resonance, of $\frac{1}{2}^{-}$ baryon octet, is obtained as in thirtyone
$\mu_{\text{val}}^{*S}\left(n^{*}\right)=-\left(\frac{2}{9}\mu_{d}^{*}+\frac{1}{9}\mu_{u}^{*}\right)\cos^{2}\theta+\left(\frac{10}{9}\mu_{d}^{*}+\frac{5}{9}\mu_{u}^{*}\right)\sin^{2}\theta\\ -\frac{8}{9}\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
$\hskip 15.6491pt\mu_{\text{sea }}^{*S}\left(n^{*}\right)=$ $\frac{a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\mu_{d}^{*}\right.\left.+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\cos^{2}\theta$
$-\frac{5a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\right.\left.\mu_{d}^{*}+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\sin^{2}\theta$
$+\frac{8a}{9}\left(1+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
$\mu_{\text{orbit }}^{*S}\left(n^{*}\right)=$ $-\left(\frac{2}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{1}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\cos^{2}\theta+\left(\frac{10}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{5}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin^{2}\theta$ (47)
$-\frac{8}{9}\left(\mu^{*}\left(d_{\uparrow}\rightarrow\right)-\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin\theta\cos\theta.$
To account for the contributions of orbital angular momentum polarizations ($\mu^{*L}_{B}$) towards the total baryonic magnetic moment, we have calculated the individual contributions from the valence quarks and quark sea as
$\mu_{B}^{*L}=\mu_{\mathrm{val}}^{*L}+\mu_{\mathrm{sea}}^{*L}.$ (48)
In this equation, $\mu_{\mathrm{val}}^{*L}$ and $\mu_{\mathrm{sea}}^{*L}$ represent the effective contributions of the valence and sea quarks towards the baryonic magnetic moment component, $\mu^{*L}_{B}$, due to the orbital angular momentum polarization. The details of the valence quark calculations has been discussed in ref. thirtyone . In the current section we will also describe the expressions used in the calculation of sea quark spin polarization in a baryon, which are obtained by substituting the third component of the orbital angular momentum of sea quarks in place of each valence quark as,
$q^{(\pm 1)}\rightarrow-T_{[q,GB]}q^{(\pm 1)}+\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}.\hskip 150.79968pt$ (49)
In eq. 49, $T_{[q,GB]}$ refers to the emission probability of a goldstone bosons (GB) from a quark $q^{(\pm 1)}$ and $\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}$ represents the transformation probability of a $q^{(\pm 1)}$ quark into two other quarks of $SU(3)$ model, given in thirtyone as
$\left|\psi\left(u^{(\pm 1)}\right)\right|^{2}=ad^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$ (50)
$\left|\psi\left(d^{(\pm 1)}\right)\right|^{2}=au^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$
$\left|\psi\left(s^{(\pm 1)}\right)\right|^{2}=a\alpha^{2}u^{(\pm 1)}+a\alpha^{2}d^{(\pm 1)}.$
The probability of transition a $q^{\pm 1}$ constituent quark into all the possible quarks ($u,d$ and $s$) along with GBs $\left(q\bar{q}^{\prime}\right)$ can be expressed as fiftyeight
$P\left(u^{\pm}\rightarrow u\bar{q}^{\prime}+q^{\prime\mp}\right)=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\mp}+ad^{\mp}+a\alpha^{2}s^{\mp},$ (51)
$P\left(d^{\pm}\rightarrow d\bar{q}^{\prime}+q^{\prime\mp}\right)=au^{\mp}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\mp}+a\alpha^{2}s^{\mp},$
$P\left(s^{\pm}\rightarrow s\bar{q}^{\prime}+q^{\prime\mp}\right)=a\alpha^{2}u^{\mp}+a\alpha^{2}d^{\mp}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\mp}.$
Using eq. 51, we can easily obtain the $T_{[q,GB]}$ for each constituent quark. The valence and sea quark contributions of the orbital angular momentum part of the magnetic moment of the baryons, can thus be calculated easily. As an example, we have presented the orbital angular momentum contribution to baryonic magnetic moment of $\frac{1}{2}^{-}$ $n^{*}$ resonances as,
$\mu^{*L}_{\mathrm{valence}}\left(n^{*}\right)=$ $\left(\frac{4}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\cos^{2}\theta-\left(\frac{1}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\sin^{2}\theta,$ (52)
$\mu^{*L}_{\mathrm{sea}}\left(n^{*}\right)=$ $\frac{2a}{9}\left[-\left(1+2\alpha^{2}\right)\mu^{*}_{d}+\left(1-\alpha^{2}\right)\mu^{*}_{u}\right.\left.+3\alpha^{2}\mu^{*}_{s}\right]\cos^{2}\theta$
$+\frac{a}{9}\left[-\left(1-\alpha^{2}\right)\mu^{*}_{d}\right.\left.+\left(1+2\alpha^{2}\right)\mu^{*}_{u}-3\alpha^{2}\mu^{*}_{s}\right]\sin^{2}\theta.$
Using similar methodology as demonstrated above, the magnetic moments of the $\frac{1}{2}^{-}$ octet baryon resonances $p^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$ and $\Xi^{*0}$ as well as the low-lying $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$, and $S_{11}^{0}(1650)$ have been calculated within the $\chi\mathrm{CQM}$ framework.
IV Results & Discussion
We now discuss numerical results on the in-medium masses (section IV.1) and magnetic moments (section IV.2) of spin $\frac{1}{2}^{-}$ octet baryon resonances. The effective baryon masses have been listed in tables 1 and 2, so as to facilitate a discussion on the mass modifications. As an outcome of the calculations discussed in the previous sections, the vacuum and in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances have been illustrated in tables 3, 4, 5 and 6. For the sake of comparison with the baryonic magnetic moments obtained from earlier studies, we have tabulated the values calculated through different models in table 7 for $\rho_{B}=0$.
IV.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
Figure 1: Effective masses of constituent quarks $u,d$ and $s$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
As the medium modifications of baryon masses is calculated using the medium modified masses of constituent quarks, we first discuss the impact of density and temperature on effective masses of quarks. In fig. 1, the in-medium masses, $m^{*}_{u},m^{*}_{d}$ and $m^{*}_{s}$, of the constituent quarks are shown as a function of density in units of nuclear saturation density, at $T=0$ MeV (on the left) and $T=100$ MeV (on the right)(where nuclear saturation density $\rho_{0}=0.16$ $\text{fm}^{-3}$). It also illustrates the in-medium modification in quark masses for different values of isospin asymmetry parameter ($i_{a}$) and strangeness fraction ($f_{s}$).
From fig. 1 it is observed that as we move from vacuum ($\rho_{B}=0$) to denser hadronic medium, the quark masses decrease monotonously. For symmetric nuclear matter at temperature $T=0$ MeV, the effective masses of light quarks decreases by 71.14% and that of $s$ quark decreases by 14.65%, at density value of $\rho_{B}=3\rho_{0}$. As observed, the effects of density on the $s$ quark is significantly lower than that on the light quarks, which is an implication to the negligible coupling of $s$ quark with $\sigma$ field fiftyfour ; sixtytwo . By introducing finite strangeness in the symmetric nuclear medium, at $f_{s}=0.3$ and $T=0$ MeV, with an increase in baryon density there is an appreciable drop in $m^{*}_{s}$, relative to that in symmetric nuclear matter, while in the light quarks only slight variations in the $m^{*}_{u}$ and $m^{*}_{d}$ values are observed. It is also observed that with the rise in temperature at $T=100$ MeV the effective masses of $u$ and $d$ quarks rises by 8.91 % at $\rho_{B}=\rho_{0}$ and the $s$ quark mass elevates by just 1.38 %. Thus an increase in the medium temperature causes a rise in the masses of constituent quarks, at the same density. The increase in isospin asymmetry of the medium leads to increase in-medium masses of constituent quarks. A study on the basis of an extended NJL model has previously demonstrated that while the in-medium masses of light quarks approach their current quark masses, the $s$ quark condensate remains large even at densities close to $\rho_{B}\approx 1$ $\text{fm}^{-3}$ xia .
In figs. 2, 3 and 4 and fig. 5 we have shown the variations in effective baryonic masses, as a function of medium density, for $\frac{1}{2}^{-}$ octet baryon resonances and $N^{*}$ resonance $S^{+(0)}_{11}(1650)$, respectively. Using eq. 25 the effective masses of $\frac{1}{2}^{-}$ baryon resonances are calculated in the isospin asymmetric strange matter, where the $E_{ispin}$ values are best fitted to procure the masses of $\frac{1}{2}^{-}$ baryon resonances in free space. A quick glance at the arguments and figures depicts, a monotonous decrease in the in-medium baryon masses as a function of density. The baryon masses in the medium fall rapidly in the lower density region up to $\rho_{B}=\rho_{0}$ and then saturate as the medium approaches large densities. As seen from fig. 2, when symmetric nuclear matter density rises upto $\rho_{B}=3\rho_{0}$, the mass of $p^{*}$ resonance drops appreciably by 33.2%. At a density of $3\rho_{0}$, with finite isospin asymmetry $i_{a}=0.3$, the mass of $p^{*}$ is 1010.03 MeV which is a rise from 1003.60 MeV at $i_{a}=0$. In fact, with higher valence $d$ quark content in a baryon, the density effects are found to be suppressed. This can be seen from the 25.27% decrease in the effective mass of $n^{*}$, as the medium density reaches $\rho_{B}=3\rho_{0}$ from vacuum. Keeping the density unchanged, say at $\rho_{B}=3\rho_{0}$, the $M^{*}_{n}$ value is found to incur a marginal increase, as a function of asymmetry of the medium, to attain 1149.18(1147.02) MeV for $i_{a}=0.3(0)$. In contrary to light quark content baryons, for symmetric nuclear matter the mass modifications incurred by $\Sigma^{*}$ resonances, as seen in fig. 3, is 18.5% and that of $\Xi^{*}$ resonances in fig. 4 is 14.5%. We can clearly state that, the density dependency on the baryons constituted by light quarks is much higher than that of hyperons, which is because $m^{*}_{u}$ and $m^{*}_{d}$ are more sensitive to density changes as compared to $m^{*}_{s}$. Additionally in hyperons, the in-medium effects are subject to the extent of $s$ quarks constituting a baryon. This is because as the number of constituent $s$ quarks increase, the in-medium effects on the constituent quark masses start having a dominating influence on the total mass of the baryon sixtytwo . In fig. 5, the masses of $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ varying as a function of medium density, have been shown. The extent of in-medium variations of the $N^{*}$ resonances is large compared to hyperons, like, in the case of $p^{*}$ and $n^{*}$ resonances. At a given value of $\rho_{B}$, with $i_{a}=0,T=0$ MeV, the impact of strangeness fraction, $f_{s}$, causes a sizable decrease in the masses of baryons. Looking up the mass of $\Xi^{*-}$ from table 1, at $\rho_{B}=3\rho_{0},f_{s}=0.3(0)$, we observe that the value decreases to 1399.14(1439.75) MeV. A cursory look at the plots of effective baryonic masses versus medium density suggests that, if the temperature of the medium increases to $T=100$ MeV then the mass modification curves will have relatively less steeper slope than that at $T=0$ MeV. For example, at $i_{a}=0,f_{s}=0(0.3)$ when the medium density is at $\rho_{B}=3\rho_{0}$, the effective mass of $\Sigma^{*+}$ being of 1300.88(1276.30) MeV, in table 1, at $T=0$ MeV increases to 1322.58(1287.69) MeV, in table 2, at $T=100$ MeV.
Owing to the strong coupling of $N^{*}$(1535) nucleon resonance with $\eta N$, there has been an earlier work that leverages $\eta$ mesic nuclei as a tool to map the reduction in-medium mass difference between nucleon $N$ and its chiral partner $N^{*}$, in chiral doublet model 2002_mass . In another study conducted using parity doublet model at finite baryonic densities of cold nuclear matter, the authors showcases the dependency of the vacuum mass of $N^{*}$ to determine the critical density and the order of the phase transition 2006_mass . For negative parity nucleons, the density effects on the in-medium baryonic mass turns out to be weaker than the impact on nucleon mass 2016_mass , which also gets reflected in fig. 2. In the context of hot hadronic matter, a study based on hadron resonance gas (HRG) sheds light on the temperature dependency of the effective masses of negative parity baryons 2018_mass . The observed temperature effects within $T=100$ MeV, have been in good agreement with the findings in current work. In the context of $\frac{1}{2}^{+}$ octet baryons, there has been studies on their in-medium mass modifications based on QMC hadron_mass , modified QMC (MQMC) mqmc , CQMF sixtytwo and also for decuplet, low-lying strange, charm, and bottom baryons in QMC framework tsushima1 ; tsushima2 .
Figure 2: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 3: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 4: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 5: Effective masses of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
IV.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
Figure 6: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 7: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 8: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 9: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 10: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 11: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 12: Effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 13: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
In this section we elaborately discuss the results of the vacuum and in-medium magnetic moments of ${\frac{1}{2}}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$, $\Xi^{*0}$ and low-lying $N^{*}$ resonances ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$, observed from figs. 6, 7, 8, 9, 10, 11, 12 and 13. For the sake of discussion, we have also included the individual contributions from valence and sea quarks to the total baryonic magnetic moment which allows us to better interpret the results.
In fig. 6, we have shown the magnetic moment of $p^{*}$ and $n^{*}$ resonances with density of the medium at temperatures $T=0$ MeV and $T=100$ MeV. The observed in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$, shows a steeper slope at lower densities of nuclear matter ($\rho_{B}\leq 2\rho_{0}$) and gradually attains saturation at much higher densities. As we trace the plot of these effective baryon magnetic moments while the medium density rises to $\rho_{B}=3\rho_{0}$, we observe that $\mu^{*}_{p^{*}}$ increases by $0.59\mu_{N}$ and $\mu^{*}_{n^{*}}$ initially approaches a minima at $-1.382\mu_{N}$ (at $\rho_{B}\approx\rho_{0}$) and then climbs up close to its vacuum value at $-1.375\mu_{N}$. A conclusive understanding of the observed effective baryonic magnetic moments of $p^{*}$ and $n^{*}$ can be provided by analyzing fig. 7, where the contributions from valence quarks, sea quarks and orbital angular momentum of the quark sea have been shown individually. For example, in case of $n^{*}$ in symmetric nuclear matter, the decrease in the valence quark contributions dominates over the increasing contributions of sea quarks and the orbital contributions of sea quarks, which leads to an overall decreasing behavior of $\mu^{*}_{n^{*}}$ below the saturation density of the medium ($\rho_{0}$). As the medium density moves past this point ($\rho_{B}\approx\rho_{0}$), owing to the dominance of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions the effective in-medium behavior of $\mu^{*}_{n^{*}}$ reverses and it approaches higher values. On a similar note, the density dependency of $\mu^{*}_{p^{*}}$ can be comprehended by taking into account the increase in $\mu^{*}_{val}$ dominating over the decrease in $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$. In symmetric strange matter at $f_{s}=0.3$, the in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ is enhanced. From fig. 6 we can infer that, the finite strangeness fraction in the medium leads to a splitting in baryonic magnetic moments of $p^{*}$ and $n^{*}$, which is visible beyond the nuclear saturation density, $\rho_{0}$. All other medium parameters affixed, non-zero isospin asymmetry ($i_{a}=0,0.3,0.5$) causes fine splitting in the $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ curves as can be seen from fig. 6, which has also been identified in their respective $\mu^{*}_{val}$, $\mu^{*}_{sea}$, $\mu^{*}_{orbit}$ contributions by referring to fig. 7. As we have included “Goldstone Boson Exchange”(GBE) effect in this study, it is important to note that the sea quark effect has an opposite sign as compared to the valence quark effect, whereas the orbital angular momentum of sea quarks has the same contribution as that from the valence quark effect (see Table 3). The sign of these contributions can be justified, if we realize that the production of the quark sea is associated with a spin flip process (see eq. 30) and hence, the sea quarks has to have an opposite effect. An earlier study highlights the impact of Goldstone Boson (GB) degrees in chiral quark model ($\chi$QM), where the magnetic moments calculated for $p^{*}$ and $n^{*}$ in vacuum, by considering GBE interaction, are 1.4 $\mu_{N}$ and $-0.9\mu_{N}$ for $\chi$QM, which are slightly lower than the corresponding values, 1.7 $\mu_{N}$ and $-1.1\mu_{N}$ as obtained from CQM framework liu . In addition, the magnetic moments procured from $\chi$QM by considering Goldstone Boson Exchange (GBE) and One Gluon Exchange (OGE), separately, gives similar magnetic moment values and are in close agreement to our vacuum state results liu . Two other studies in the literature, with quasibound state approach quasibound and chiral unitary model chiral_unitary have previously calculated for magnetic moments of these two resonances and could be compared from table 7. Table 7 facilitates a comparitive study of the magnetic moments of $p^{*}$ and $n^{*}$ resonances obtained from various models.
Figure 8 depicts the in-medium modifications of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$, as a function of $\rho_{B}/\rho_{0}$ and the behavior of the individual contributions have been plotted in fig. 9. In nuclear medium, the magnetic moments of $\Sigma^{*+}$ initially increases up to $\rho_{B}=2\rho_{0}$ and then saturates at higher densities of the medium. By taking into account the individual contributions from fig. 9, we see that $\mu^{*}_{val}$ increases with density and it has a dominating effect over the cumulative depletion of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions. In case of $\Sigma^{*-}$ (at $i_{a}=0,f_{s}=0$), the effective magnetic moment $\mu^{*}_{\Sigma^{*-}}$ decreases and gradually approaches a saturation at very high densities of $\rho_{B}\approx 5\rho_{0}$. Although the in-medium behavior of $\mu^{*}_{val}$ is accompanied by $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$, which imparts opposite contributions, but the valence quarks have a predominant role in determining the nature of $\mu^{*}_{\Sigma^{*-}}$. These opposite contributions due to sea quarks and the orbital motion of sea quarks at higher density regime, evidently explains the stagnant in-medium nature of $\mu^{*}_{\Sigma^{*-}}$ at large medium densities. Further, the density dependent medium modification of $\mu^{*}_{\Sigma^{*0}}$, in symmetric nuclear matter, is much alike the in-medium nature of $\mu^{*}_{\Sigma^{*+}}$ except for a slightly decreasing slope at densities above $1.5\rho_{0}$. The dominance of $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions over the $\mu^{*}_{sea}$ contribution, justifies the in-medium curve of $\mu^{*}_{\Sigma^{*0}}$ in a nuclear medium. From a point of comparison if you see Table 7, the magnetic moments of $\mu^{*}_{\Sigma^{*+}}$ (1.647 $\mu_{N}$), $\mu^{*}_{\Sigma^{*-}}$ $(-0.754\mu_{N})$ and $\mu^{*}_{\Sigma^{*0}}$ (0.446 $\mu_{N}$) in free space are found to be in agreement with the NCQM results 1.814 $\mu_{N}$, $-0.689\mu_{N}$ and 0.820 $\mu_{N}$, respectively. As seen from Tables 3 and 5, for $f_{s}=0(0.3)$ at $\rho_{B}=3\rho_{0}$, $\mu^{*}_{\Sigma^{*+}}$ and $\mu^{*}_{\Sigma^{*-}}$ values attain 1.697(1.723) $\mu_{N}$ and $-0.770(-0.779)\mu_{N}$. This suggests that as we introduce finite strangeness fraction to the non-strange medium, there will be an enhancement in the nature of in-medium variations of baryonic magnetic moments of $\Sigma^{*+}$, $\Sigma^{*0}$ and $\Sigma^{*-}$. For finite $f_{s}$, the magnetic moments of $\Sigma^{*+}$ and $\Sigma^{*0}$ surges while that of $\Sigma^{*-}$ approaches lower values, instead of attaining a saturation at large densities. For an instance, the in-medium values of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ (in units of $\mu_{N}$) at $i_{a}=0(0.5)$, when referred to Tables 3 and 4, are $1.697(1.698),-0.770(-0.763)and0.463(0.472)$ in a medium density of $3\rho_{0}$. This and a glance into fig. 8 makes it apparent that the impact of non-zero $i_{a}$, a fine splitting in effective baryonic magnetic moments, is more visible for $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ as compared to $\mu^{*}_{\Sigma^{*+}}$. Such $i_{a}$ dependent fine shifts in the baryonic magnetic moment have been previously found in works, in the present formalism, dedicated to octet baryons fiftythree and decuplet baryons fiftyfour .
We find it interesting to note that for $\Xi^{*}$ baryons the density effects on $\mu^{*}_{\Xi^{*-}}$ and $\mu^{*}_{\Xi^{*0}}$ values show impressive similarity, owing to the valence quark spin polarizations and corresponding orbital moments of two $s$ quarks that suppresses the impact of light $u$ and $d$ quarks in $\Xi^{*-}$ and $\Xi^{*0}$, respectively sixtytwo . The baryonic magnetic moment of $\Xi^{*-}$, as seen from fig. 10, rises steeply with increase in the relative density ($\rho_{B}/\rho_{0}$) of the medium. Figure 11 clearly illustrates that for $\Xi^{*-}$, the cumulative increasing behavior of $\mu^{*}_{val}$ and $\mu^{*}_{sea}$ contributions overshadows the decreasing orbital contributions from sea quarks ($\mu^{*}_{orbit}$). In a similar manner, a quantitative analysis of fig. 10 and fig. 11 suggests that the density dependent increase in $\mu^{*}_{\Xi^{*0}}$ is well justified by the contributions coming from $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$, which suppresses the contributions coming from $\mu^{*}_{sea}$. Much alike other baryons, we have observed a correlation between finite isospin asymmetry of the medium and a distinguishable splitting in the baryonic magnetic moments in case of $\Xi^{*}$ resonances. Now in order to understand the temperature effects on the in-medium baryonic magnetic moments, we shall lay emphasis on the case of $\mu^{*}_{\Xi^{*-}}$. A close look at fig. 10, shows that for symmetric nuclear matter at medium densities of $\rho_{0}$ and $3\rho_{0}$, at $T=0(100)$ MeV the $\mu^{*}_{\Xi^{*-}}$ values obtained, in $\mu_{N}$ units, are 0.313(0.286) and 0.179(0.163), respectively. This reveals the in-medium nature of $\mu^{*}_{\Xi^{*-}}$, an increasing curve as a function of medium density, is subject to slight reduction in its value as the temperature rises to $T=100$ MeV. In addition, the vacuum state magnetic moments of $\Xi^{*-}$ resonances are found to be much lower than those procured from NCQM (see Table 7).
From fig. 12, we observe that the magnetic moments of ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$ increases, overall as a function of density of the hadronic matter $\rho_{B}$ (in $\rho_{0}$ units). As from fig. 13 it is seen that, the $\mu^{*}_{sea}$ exercises control over the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions, which explains the medium modifications of the magnetic moment of ${S^{+}_{11}}(1650)$ in lower density regime (i.e. up till $\rho_{B}=2\rho_{0}$). However, as the nuclear medium grows denser the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions demonstrates a dominating influence on the total effective magnetic moment of ${S^{+}_{11}}(1650)$. From fig. 13 it is also observed that for the $N^{*}$ resonance ${S^{0}_{11}}(1650)$, the value of $\mu_{{S^{0}_{11}}(1650)}$ receives major contributions, monotonously, due to the dominant valence quark contribution ($\mu^{*}_{val}$) decreasing with rise in density. For a comparative study based on different models, similar justifications as discussed in the case of $p^{*}$ and $n^{*}$ resonances suffices and the results have been tabulated in Table 7.
A study based on QMC as well as MQMC, for octet baryons, shows that we can expect sizable influence on effective baryonic moments due to temperature above $T=150$ MeV at lower medium densities mqmc . Speaking of correlations, the authors in ref. seventyfour have discussed a strong correlation of the baryonic magnetic moment with that of the bag radius of a baryon, as the two quantities have a similar behavior as density dependent functions. Since this is the first work on medium modifications of $\frac{1}{2}^{-}$ baryon resonances, so referring to previous works on in-medium density effects on magnetic moments of baryons using the current formalism sixtytwo ; fiftyfour and QMC tsushima2 , might facilitate the purpose of a comparative study for future work on the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances.
V Summary
To summarize, in the present work we have investigated the modifications induced on the masses as well as the magnetic moments of $J_{p}={\frac{1}{2}}^{-}$ baryon resonances in isospin asymmetric strange hadronic matter. In light of chiral $SU(3)$ quark mean field model the in-medium effects are introduced and we computed the effective baryon masses by using the masses of constituent quarks. The mass adjusted quark magnetic moments are then calculated, to include relativistic quark confinement, by using the baryon masses as inputs. In hadronic medium the in-medium masses of lights quarks, $u$ $d$, vary significantly in comparison to the observed changes in medium modified masses of $s$ quark. In the $\chi$CQM formalism, we then computed the baryonic magnetic moments from the contributions due to valence quarks, sea quarks and orbital angular momentum due of sea quarks. As derived from the results is that, the medium modifications are more prominent for the baryons comprised of light $u$ or $d$ quarks, than for hyperons. For strange matter, the in-medium constituent quark masses, effective baryonic masses as well as the effective baryonic magnetic moments suffer appreciable enhancement in their in-medium behaviour. By carrying out this study at finite temperatures and finite $i_{a}$ values of the hadronic medium, we have noted a fine shift in the in-medium baryonic magnetic moments, although the temperature effects have been negligible below $T=150$ MeV.
Baryon Vacuum Mass $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$f_{s}=0$ $f_{s}=0.3$ $f_{s}=0$ $f_{s}=0.3$
(MeV) $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$
$M_{S^{+}_{11}(1650)}^{*}$ 1650.0 $1439.01$ $1442.02$ $1435.09$ $1440.78$ $1263.40$ $1268.27$ $1254.90$ $1260.31$
$M_{S^{0}_{11}(1650)}^{*}$ 1650.0 $1439.01$ $1439.98$ $1435.09$ $1438.69$ $1263.40$ $1265.55$ $1254.90$ $1257.48$
$M_{p^{*}}^{*}$ 1535.0 $1241.14$ $1245.05$ $1225.47$ $1233.20$ $1003.60$ $1010.03$ $953.84$ $962.62$
$M_{n^{*}}^{*}$ 1535.0 $1323.31$ $1324.28$ $1319.37$ $1322.98$ $1147.02$ $1149.18$ $1138.50$ $1141.08$
$M_{\Sigma^{*+}}^{*}$ 1620.0 $1442.84$ $1446.59$ $1434.69$ $1440.74$ $1300.88$ $1306.54$ $1276.30$ $1283.18$
$M_{\Sigma^{*-}}^{*}$ 1620.0 $1442.84$ $1442.48$ $1434.69$ $1436.54$ $1300.88$ $1301.06$ $1276.30$ $1277.49$
$M_{\Sigma^{*0}}^{*}$ 1620.0 $1442.84$ $1444.53$ $1434.69$ $1438.64$ $1300.88$ $1303.80$ $1276.30$ $1280.34$
$M_{\Xi^{*-}}^{*}$ 1690.0 $1547.34$ $1547.70$ $1534.99$ $1537.17$ $1439.75$ $1440.69$ $1399.14$ $1401.79$
$M_{\Xi^{*0}}^{*}$ 1690.0 $1547.34$ $1549.75$ $1534.99$ $1539.27$ $1439.75$ $1443.43$ $1399.14$ $1404.64$
[…]
Baryon Vacuum Mass $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$f_{s}=0$ $f_{s}=0.3$ $f_{s}=0$ $f_{s}=0.3$
(MeV) $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$
$M_{S^{+}_{11}(1650)}^{*}$ 1650.0 $1469.66$ $1471.49$ $1454.68$ $1455.95$ $1289.10$ $1292.78$ $1270.06$ $1272.66$
$M_{S^{0}_{11}(1650)}^{*}$ 1650.0 $1469.66$ $1469.70$ $1454.68$ $1453.87$ $1289.10$ $1289.73$ $1270.06$ $1269.38$
$M_{p^{*}}^{*}$ 1535.0 $1283.87$ $1286.17$ $1251.81$ $1253.30$ $1040.65$ $1045.23$ $972.85$ $976.01$
$M_{n^{*}}^{*}$ 1535.0 $1354.07$ $1354.10$ $1339.04$ $1338.22$ $1172.82$ $1173.45$ $1153.71$ $1153.02$
$M_{\Sigma^{*+}}^{*}$ 1620.0 $1468.45$ $1471.06$ $1450.57$ $1452.89$ $1322.58$ $1327.39$ $1287.69$ $1291.80$
$M_{\Sigma^{*-}}^{*}$ 1620.0 $1468.45$ $1467.46$ $1450.57$ $1448.72$ $1322.58$ $1321.26$ $1287.69$ $1285.20$
$M_{\Sigma^{*0}}^{*}$ 1620.0 $1468.45$ $1469.26$ $1450.57$ $1450.81$ $1322.58$ $1324.33$ $1287.69$ $1288.50$
$M_{\Xi^{*-}}^{*}$ 1690.0 $1567.79$ $1567.58$ $1547.08$ $1546.28$ $1457.33$ $1457.12$ $1406.70$ $1405.71$
$M_{\Xi^{*0}}^{*}$ 1690.0 $1567.79$ $1569.38$ $1547.08$ $1548.36$ $1457.33$ $1460.19$ $1406.70$ $1409.02$
[…]
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.161 -0.200 -0.051 -0.090 0.184 -0.236 0.061 0.008
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.065 -0.286 0.125 0.904 1.213 -0.334 0.064 0.943
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.299 -0.392 0.207 2.114 2.843 -0.483 0.016 2.376
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.464 0.263 -0.181 -1.382 -1.689 0.305 0.007 -1.376
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.935 -0.382 0.132 1.684 2.086 -0.419 0.030 1.697
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.821 0.153 -0.091 -0.760 -0.971 0.174 0.026 -0.770
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.556 -0.114 0.020 0.461 0.557 -0.122 0.028 0.463
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.106 0.231 0.052 0.177 0.093 0.233 -0.017 0.309
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.833 -0.282 -0.084 -1.200 -0.687 -0.318 -0.019 -1.026
Table 3: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric nuclear matter ($i_{a}=0$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.160 -0.199 -0.047 -0.085 0.182 -0.233 0.064 0.013
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.062 -0.285 0.134 0.911 1.205 -0.331 0.079 0.953
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.280 -0.389 0.213 2.104 2.792 -0.475 0.032 2.349
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.459 0.263 -0.185 -1.382 -1.676 0.303 -0.003 -1.376
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.927 -0.380 0.137 1.684 2.072 -0.416 0.042 1.698
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.818 0.152 -0.092 -0.757 -0.960 0.173 0.024 -0.763
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.557 -0.114 0.023 0.466 0.561 -0.122 0.033 0.472
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.116 0.231 0.052 0.168 0.065 0.234 -0.015 0.284
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.840 -0.281 -0.087 -1.209 -0.710 -0.314 -0.026 -1.051
Table 4: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric nuclear matter ($i_{a}=0.5$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.162 -0.200 -0.049 -0.088 0.185 -0.238 0.068 0.015
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.068 -0.287 0.124 0.904 1.221 -0.337 0.062 0.946
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.328 -0.397 0.205 2.136 2.991 -0.509 0.005 2.487
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.468 0.264 -0.178 -1.382 -1.701 0.308 0.018 -1.375
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.946 -0.384 0.130 1.692 2.125 -0.427 0.026 1.723
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.826 0.153 -0.089 -0.762 -0.991 0.177 0.034 -0.779
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.560 -0.115 0.020 0.464 0.567 -0.125 0.030 0.472
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.108 0.233 0.051 0.176 0.100 0.239 -0.021 0.319
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.841 -0.285 -0.083 -1.209 -0.703 -0.328 -0.016 -1.048
Table 5: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric hyperonic matter ($i_{a}=0$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.160 -0.198 -0.047 -0.085 0.183 -0.235 0.074 0.022
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.061 -0.284 0.135 0.911 1.216 -0.334 0.075 0.957
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.295 -0.392 0.216 2.119 2.944 -0.501 0.016 2.459
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.458 0.262 -0.187 -1.382 -1.694 0.306 0.012 -1.375
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.933 -0.381 0.138 1.690 2.112 -0.424 0.035 1.723
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.818 0.152 -0.093 -0.758 -0.982 0.176 0.034 -0.771
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.560 -0.115 0.023 0.468 0.570 -0.124 0.034 0.480
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.122 0.233 0.053 0.163 0.076 0.240 -0.021 0.295
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.851 -0.282 -0.088 -1.222 -0.722 -0.324 -0.022 -1.069
Table 6: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric hyperonic matter ($i_{a}=0.5$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
Baryons $\chi$CQM $\chi$QMliu CQMliu NCQMthirtyone Quasiboundquasibound CUchiral_unitary
$\mu_{\text{val}}$ $\mu_{\text{sea}}$ $\mu_{\text{orbit}}$ $\mu_{B}$ $\mu_{B}$ (GBE) $\mu_{B}$ (OGE) $\mu_{B}$ (GBE) $\mu_{B}$ $\mu_{B}$ $\mu_{B}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.1 0.0 0.3 0.106 - -
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 0.6 0.7 0.8 0.951 - -
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 1.4 1.6 1.7 1.894 1.86 1.13
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -0.9 -1.0 -1.1 -1.284 -0.56 -0.248
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 - - - 1.814 - -
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 - - - -0.689 - -
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 - - - 0.820 - -
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 - - - -0.315 - -
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 - - - -0.990 - -
Table 7: The above table allows a comparative study of magnetic moments of $\frac{1}{2}^{-}$ baryon resonances, computed using different models in the free space ($\rho_{B}=0$).
VI APPENDIX
By minimizing the thermopotential in eq. 27, we derive the equations of motion for scalar fields $\sigma$, $\zeta$, the dilaton field $\chi$, scalar iso-vector field $\delta$ and the vector fields $\omega,\rho,\phi$ in view of eq. 29 as in fiftyfour
[…]
$\,+\,\frac{\chi^{2}}{\chi_{0}^{2}}m_{\pi}^{2}f_{\pi}-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\omega}\omega^{2}\frac{\partial m_{\omega}}{\partial\sigma}\,-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\rho}\rho^{2}\frac{\partial m_{\rho}}{\partial\sigma}\,+\,\sum_{i}\frac{\partial M_{i}^{\ast}}{\partial\sigma}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (53)
[…]
$+\,\frac{\chi^{2}}{\chi_{0}^{2}}\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)+\,\sum_{i=\Lambda,\Sigma^{\pm,0},\Xi^{-,0}}\frac{\partial M_{i}^{\ast}}{\partial\zeta}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}=0,$ (54)
$\frac{\partial\Omega}{\partial\chi}=k_{0}\chi\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)-k_{3}\left(\sigma^{2}-\delta^{2}\right)\zeta+\frac{2\chi}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)\zeta\right]$
$-\frac{\chi}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}\right)+\left(4k_{4}+1+{\rm ln}\frac{\chi^{4}}{\chi_{0}^{4}}-\frac{4\xi}{3}{\rm ln}\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\right)\right)\chi^{3}=0,$ (55)
[…]
$-\,\sum_{i}g_{\delta i}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (56)
$\frac{\partial\Omega}{\partial\omega}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\omega}^{2}\omega+4g_{4}\omega^{3}+12g_{4}\omega\rho^{2}-\sum_{i}g_{\omega i}\rho_{i}=0,$ (57)
$\frac{\partial\Omega}{\partial\rho}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\rho}^{2}\rho+4g_{4}\rho^{3}+12g_{4}\omega^{2}\rho-\sum_{i}{g_{\rho i}}\rho_{i}=0,$ (58)
and
$\frac{\partial\Omega}{\partial\phi}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\phi}^{2}\phi+8g_{4}\phi^{3}-\sum_{i}g_{\phi\,i}\rho_{i}=0.$ (59)
In eqs. 57, 58 and 59 $\rho_{i}$ represents the vector density of baryons and is given by,
$\rho_{i}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\Big{(}f_{i}(k)-\bar{f}_{i}(k)\Big{)},$ (60)
and in sections VI, VI and VI $\rho_{i}^{s}$ stands for the scalar density of baryons, given as
$\rho_{i}^{s}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{M_{i}^{*}}{E^{\ast}_{i}(k)}\Big{(}f_{i}(k)+\bar{f}_{i}(k)\Big{)}.$ (61)
Here, $f_{i}(k)$ and $\bar{f}_{i}(k)$ appearing in eqs. 60 and 61 represent the Fermi distribution functions for fermions and anti-fermions, at finite temperature and are expressed as
$f_{i}(k)$ $=$ $\frac{1}{1+\exp\left[(E^{*}_{i}(k)-\nu^{*}_{i})/k_{B}T\right]}~{}~{}~{}\text{and}~{}~{}~{}\bar{f}_{i}(k)=\frac{1}{1+\exp\left[(E^{*}_{i}(k)+\nu^{*}_{i})/k_{B}T\right]}~{}.$ (62)